# A Note on Output Length of One-Way State Generators and EFIs

### v3

@@ 1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22b, MY22a], and EFIs [BCQ23] are quantum analogue of PRGs, OWFs, and EFIDs [Gol90], respectively. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22b, AQY22, BCQ23, Yan22].
OWSGs are a quantum analog of OWFs.111 There are a pure state output version [MY22b] and mixed state output version [MY22a] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^2] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
@@
An EFI is a pair of efficiently generatable states that are statistically far but computationally indistinguishable. While relationships between EFIs and other primitives have been actively studied [MY22a, BCQ23], to our knowledge, there is no existing work that studies how many qubits are needed for EFIs.
@@ 1.1 Our Results
In this regime, the idea for the proofs of Theorems 1.1 and 1.2 does not work since Welch bound only gives an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. Thus, we rely on a different idea. Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY23]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^6] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.10 and 2.12) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Finally, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY22a]. We show the following lower bound.
@@ 1.2 Discussion
    Quantum-public-key digital signatures: OWSGs are equivalent to (bounded poly-time secure) digital signatures with quantum public keys [MY22a]. Hence we have the following consequences. First, from Theorem 1.1, the length of the quantum public key of digital signatures cannot be $O(\log\lambda)$. Second, from Theorem 4.1, assuming OWFs, there exist one-time secure digital signatures with $m$-qubit quantum public keys for any $m=\omega(\log(\lambda))$. This can be regarded as a quantum advantage because it is unlikely that we can construct classical one-time secure digital signatures with similar public key length from (polynomially secure) OWFs.
@@
    Private-key quantum money: Private-key quantum money schemes with pure banknotes imply OWSGs [MY22a]. Hence, from Theorem 1.1, the length of pure quantum banknotes cannot be $O(\log\lambda)$.
@@
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY23] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proofs of Theorems 1.3 and 1.4.
@@ 2.1 Basic Cryptographic Primitives
Definition 2.4 (One-way state generators (OWSGs)  [MY22a]).
@@ 2.2 Useful Lemmata
Lemma 2.10 (Quantum state tomography [ALY23, Corollary 2.16]).
@@
[ALY23, Corollary 2.16] does not mention that the output matrix is Hermitian. This fact can be checked by following the proof; the starting point is the folklore tomography (see [Low21, Section 1.5.3]) that approximates $\rho=\frac{1}{d}\sum\mathrm{Tr}(P_{i}\rho)P_{i}$ for Pauli matrices $P_{i}$ by $N=\sum\mu_{i}P_{i}$ for $\mu_{i}\approx\frac{1}{d}\mathrm{Tr}(P_{i}\rho)$. The matrix $N$ is Hermitian, and the above lemma, which is a trace-distance version of [AGQY22, Corollary 7.6], only chooses a nice $M$ among many independent $N$’s, so that $M$ is still Hermitian.
@@
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [BDF+11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.13. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regime which we consider in this paper.
@@ 3.1 Lower Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
where Equation 14 follows from the triangle inequality for trace distance, Equation 16 follows from the concavity of the function $\sqrt{1-x}$, and Equation 17 follows from Corollary 2.9. This completes the proof of Equation 11.
@@ 5.1 Lower Bound
We remark that $\mathcal{A}$ may not be QPT since there may not exist a polynomial size (or even finite size) quantum circuit that exactly implements the projective measurement $\{\Pi,I-\Pi\}$. However, we observe that we can approximate $\mathcal{A}$ in QPT with an exponentially small error as follows. First, tomography in Step 3 runs in QPT since $t=144\lambda 2^{4\ell}/\delta^{2}={\mathrm{poly}}(\lambda)$ and the running time of $\mathsf{Tomography}$ is polynomial in $1/\delta,d=2^{\ell}$ and $\lambda$ as required in Lemma 2.10. In Step 4, by a numerical implementation of a standard diagonalization algorithm, we can obtain a description of a projector $\widetilde{\Pi}$ (which consists of all the entries of the matrix $\widetilde{\Pi}$) such that $\|\Pi-\widetilde{\Pi}\|_{op}\leq 2^{-\Omega(\lambda)}$ in QPT where $\|\cdot\|_{op}$ means the operator norm. It is known that any $\ell$-qubit unitary can be implemented with an error $\epsilon$ (in terms of the operator norm) by a ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$-quantum circuit [NC10, Sec. 4.5]. By inspecting the proof, we can see that we can compute the description of the approximating quantum circuit in (classical) time ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$ given the description of the unitary. By setting $\epsilon:=2^{-\Omega(\lambda)}$, the above implies that we can compute a description of a quantum circuit that implements projective measurement $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ such that $\|\widetilde{\Pi}-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$. By the triangle inequality, $\|\Pi-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$, and thus the output distribution changes only by $2^{-\Omega(\lambda)}$ even if we apply $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ instead of $\{\Pi,I-\Pi\}$. This means that $\mathcal{A}$ can be approximated with an exponentially small error in QPT.
@@ 5.2 Upper Bound
Acknowledgements. This work was initiated while the authors were joining NII Shonan Meeting No.198 “New Directions in Provable Quantum Advantages”. We thank Ryu Hayakawa and Fermi Ma for useful conversations. We thank anonymous reviewers of CRYPTO 2024 and TQC 2024 for their valuable comments, especially for pointing out that injectivity is not needed in Theorem 4.7. MH is supported by a KIAS Individual Grant QP089801. TM is supported by JST CREST JPMJCR23I3, JST Moonshot R&D JPMJMS2061-5-1-1, JST FOREST, MEXT QLEAP, the Grant-in Aid for Transformative Research Areas (A) 21H05183, and the Grant-in-Aid for Scientific Research (A) No.22H00522.
@@ A.3 Upper Bound
which implies statistical binding (e.g., see [MY22b]). The number of qubits of the commitment register is $2n$. If $m$ is even, $2n=m$ and if $m$ is odd, $12n=m-$ 1, so we can simply append one-qubit to the commitment register to make it $m$-qubit.

### v4

@@ 1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22, MY24], and EFIs [BCQ23] are quantum analogue of PRGs, OWFs, and EFIDs [Gol90], respectively. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22, AQY22, BCQ23, Yan22].
OWSGs are a quantum analog of OWFs.111 There are a pure state output version [MY22] and mixed state output version [MY24] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^2] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
@@
An EFI is a pair of efficiently generatable states that are statistically far but computationally indistinguishable. While relationships between EFIs and other primitives have been actively studied [MY24, BCQ23], to our knowledge, there is no existing work that studies how many qubits are needed for EFIs.
@@ 1.1 Our Results
In this regime, the idea for the proofs of Theorems 1.1 and 1.2 does not work since Welch bound only gives an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. Thus, we rely on a different idea. Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY24]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^6] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.10 and 2.12) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Finally, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY24]. We show the following lower bound.
@@ 1.2 Discussion
    Quantum-public-key digital signatures: OWSGs are equivalent to (bounded poly-time secure) digital signatures with quantum public keys [MY24]. Hence we have the following consequences. First, from Theorem 1.1, the length of the quantum public key of digital signatures cannot be $O(\log\lambda)$. Second, from Theorem 4.1, assuming OWFs, there exist one-time secure digital signatures with $m$-qubit quantum public keys for any $m=\omega(\log(\lambda))$. This can be regarded as a quantum advantage because it is unlikely that we can construct classical one-time secure digital signatures with similar public key length from (polynomially secure) OWFs.
@@
    Private-key quantum money: Private-key quantum money schemes with pure banknotes imply OWSGs [MY24]. Hence, from Theorem 1.1, the length of pure quantum banknotes cannot be $O(\log\lambda)$.
@@
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY24] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proofs of Theorems 1.3 and 1.4.
@@ 2.1 Basic Cryptographic Primitives
Definition 2.4 (One-way state generators (OWSGs)  [MY24]).
@@ 2.2 Useful Lemmata
Lemma 2.10 (Quantum state tomography [ALY24, Corollary 2.16]).
@@
[ALY24, Corollary 2.16] does not mention that the output matrix is Hermitian. This fact can be checked by following the proof; the starting point is the folklore tomography (see [Low21, Section 1.5.3]) that approximates $\rho=\frac{1}{d}\sum\mathrm{Tr}(P_{i}\rho)P_{i}$ for Pauli matrices $P_{i}$ by $N=\sum\mu_{i}P_{i}$ for $\mu_{i}\approx\frac{1}{d}\mathrm{Tr}(P_{i}\rho)$. The matrix $N$ is Hermitian, and the above lemma, which is a trace-distance version of [AGQY22, Corollary 7.6], only chooses a nice $M$ among many independent $N$’s, so that $M$ is still Hermitian.
@@
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [BDF+11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.13. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regimes which we consider in this paper.
@@ 3.1 Lower Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
where LABEL:{eq:sum_is_one} follows from $\sum_{i}p_{k,i}=\sum_{j}p_{k^{\prime}.j}=1$ for each $k.k^{\prime}$, Equation 14 follows from the triangle inequality for trace distance, Equation 16 follows from the concavity of the function $\sqrt{1-x}$, and Equation 17 follows from Corollary 2.9. This completes the proof of Equation 11.
@@ 5.1 Lower Bound
We remark that $\mathcal{A}$ may not be QPT since there may not exist a polynomial size (or even finite size) quantum circuit that exactly implements the projective measurement $\{\Pi,I-\Pi\}$. However, we observe that we can approximate $\mathcal{A}$ in QPT with an exponentially small error as follows. First, tomography in Step 3 runs in QPT since $t=144\lambda 2^{4\ell}/\delta^{2}={\mathrm{poly}}(\lambda)$ and the running time of $\mathsf{Tomography}$ is polynomial in $1/\delta,d=2^{\ell}$ and $\lambda$ as required in Lemma 2.10. In Step 4, by a numerical implementation of a standard diagonalization algorithm, we can obtain a description of a projector $\widetilde{\Pi}$ (which consists of all the entries of the matrix $\widetilde{\Pi}$) such that $\|\Pi-\widetilde{\Pi}\|_{op}\leq 2^{-\Omega(\lambda)}$ in QPT where $\|\cdot\|_{op}$ means the operator norm. It is known that any $\ell$-qubit unitary can be implemented with an error $\epsilon$ (in terms of the operator norm) by a ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$-size quantum circuit. In fact, [NC10, Sec. 4.5] gives a constructive way to compute the description of the approximating quantum circuit in (classical) time ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$ given the description of the unitary. By setting $\epsilon:=2^{-\Omega(\lambda)}$, the above implies that we can efficiently compute a description of a quantum circuit that implements projective measurement $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ such that $\|\widetilde{\Pi}-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$. By the triangle inequality, $\|\Pi-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$, and thus the output distribution changes only by $2^{-\Omega(\lambda)}$ even if we apply $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ instead of $\{\Pi,I-\Pi\}$. This means that $\mathcal{A}$ can be approximated with an exponentially small error in QPT.
@@ 5.2 Upper Bound
Acknowledgements. This work was initiated while the authors were joining NII Shonan Meeting No.198 “New Directions in Provable Quantum Advantages”. We thank Ryu Hayakawa and Fermi Ma for useful conversations. We thank anonymous reviewers of CRYPTO 2024 and TQC 2024 for their valuable comments, especially for pointing out that injectivity is not needed in Theorem 4.7. MH is supported by a KIAS Individual Grant QP089802. TM is supported by JST CREST JPMJCR23I3, JST Moonshot R&D JPMJMS2061-5-1-1, JST FOREST, MEXT QLEAP, the Grant-in Aid for Transformative Research Areas (A) 21H05183, and the Grant-in-Aid for Scientific Research (A) No.22H00522.
@@ A.3 Upper Bound
which implies statistical binding (e.g., see [MY22]). The number of qubits of the commitment register is $2n$. If $m$ is even, $2n=m$ and if $m$ is odd, $12n=m-$ 1, so we can simply append one-qubit to the commitment register to make it $m$-qubit.